# Gyral-Sulcal-Net: An Integrated Network Representation of Brain Folding Patterns

### v1

@@ Abstract
Our brain functions as a complex communication network, and studying it from a network perspective offers valuable insights into its organizational principles and links to cognitive functions and brain disorders. However, most current network studies typically use brain regions as nodes, often overlooking the intricate folding patterns of finer-scale anatomical landmarks within these regions. In this study, we introduce a novel approach to integrate the brain’s two primary folding patterns – gyri and sulci – into a unified network termed the Gyral-Sulcal-Net (GS-Net), in which three different types of finer-scale landmarks have been successfully identified. We evaluated the proposed GS-Net across multiple datasets, comprising over 1,600 brains, spanning different age groups (from 34 gestational weeks to elderly adults) and cohorts (healthy brains and those with pathological conditions). The experimental results demonstrate that the GS-Net can effectively integrate and represent diverse cortical folding patterns from a network perspective. More importantly, this approach offers a promising way for integrating different folding patterns into a unified anatomical brain network, alongside structural and functional networks, providing a comprehensive framework for studying brain networks. (The GS-Net toolbox will be released soon.)
@@ 1 Introduction
Brain network study starts with the identification of network nodes as interacting units and their interconnections as edges. In neuroimaging studies, various parcellation methods [ARSLAN20185], including those based on anatomical landmarks (e.g. AAL), cytoarchitectonic information (e.g. Brodmann areas), and connectivity patterns, are employed to subdivide the brain into a set of brain regions, which are utilized as network nodes. Advanced imaging techniques, such as Magnetic Resonance Imaging (MRI), are employed to construct macro-scale networks based on these ROIs, including structural network inferred from diffusion MRI (dMRI) and functional network estimated using resting-state functional MRI (rs-fMRI) [doi:10.1073/pnas.0811168106, https://doi.org/10.1002/hbm.22933]. Structural network allows estimation of the physical connections, while the functional network elucidates putative functional connections. Brain network studies have made significant breakthroughs in understanding the fundamental organizational principles of normal brain [ZHANG2022102463] and brain diseases [Bassett9239]. However, current research still lacks sufficient investigation into two critical aspects. Firstly, most existing brain networks utilize brain regions as nodes, which often cover large areas at a relatively coarse scale. The folding pattern of finer-grained anatomical landmarks within these brain regions tends to be overlooked, such as sulci fundi (deepest loci of sulci) and gyri peaks (highest loci in gyri) [Thecerebralsulciandgyri]. Accumulating evidence suggest that these landmarks confer specific structural, functional, and cognitive patterns [10.1093/cercor/bhac537] and are under a stronger genetic or heritability influence than other cortical regions [10.1093/cercor/bhac537, IM2019881]. Therefore, establishing a finer-scale brain network based on these landmarks is significant and would serve as a valuable complement to the region-based brain network. Secondly, existing studies primarily employ structural and functional connectivity to construct brain networks, and anatomical folding patterns cannot be effectively integrated into such frameworks. Although certain studies have investigated the unique structural-functional patterns of specific folding patterns and their changes in disease development [KIKINIS19947, 10.3389/fpsyt.2022.1033918], these methods often independently analyze particular folding patterns. They fail to effectively integrate different folding patterns within a unified brain network, alongside the brain structural and functional networks, to comprehensively study the brain from a network perspective. To address these problems, in this study, we introduced a novel framework to integrate the brain’s two primary folding patterns – gyri and sulci, into a unified network called Gyral-Sulcal-Net (GS-Net), while also automatically identifying three important landmarks – gyri conjunctions, sulci conjunctions, and sulci-gyri conjunctions. We evaluated the proposed GS-Net using multiple datasets comprising 1,600+ brains. These datasets cover various age groups ranging from 34 gestational weeks to elderly adults, and different cohorts including healthy brains and brains with pathological conditions. The generated GS-Net can accurately describe the folding pattern of the brain from a network perspective.
@@ 2.1 Method Overview
Fig. 1: The pipeline of proposed GS-Net framework.
@@ 2.4 Tree Marching and Trimming
After segmenting the gyri crests and sulci valleys, we employed tree marching and trimming methods (Fig. 1(c)) to construct networks upon gyri crests and sulci valley. Taking the gyri crest regions as an example, the tree marching process begins with the vertex at the highest gyri altitude and progressively connects other vertices in descending order until all vertices within the gyri crests are linked. During this process, all vertices in the crest area are connected, resulting in some redundant branches extending from the skeleton to the boundary. We then perform trimming to eliminate these extraneous branches. Since the remaining gyri crests form narrow regions, the redundant branches that extend beyond the main structure are relatively short and can be easily identified. Following the trimming, the main skeleton is integrated into a network that delineates the primary morphology of the gyri regions, which we refer to as GyralNet. Similarly, SulcalNet is constructed for the sulci valley regions.
@@ 2.5 Integration and Landmark Identification
GS-Net contains three distinct types of nodes, each playing a key role in brain networks. The first type includes conjunctions of multiple edges in GyralNet, representing intersections of various gyri on the cerebral cortex. These nodes have been highlighted in the literature as regions with thicker cortices, higher fiber density, and significant variations in structural and functional connectivity [LI20101202, 10.1093/cercor/bhx227]. Based on their connections to free sulci, we further classify these nodes as GC (Gyri Conjunction) and SGC (Sulci-Gyri Conjunction) in this study. Additionally, the conjunctions of multiple edges in SulcalNet show unique structural, functional, and genetic characteristics. Existing study suggests these nodes exhibit spatial consistency across individuals during development [Thecerebralsulciandgyri]and are subject to stronger genetic or heritability influences compared to other cortical regions [10.1093/cercor/bhac537]. These nodes are referred to as SC (Sulci conjunctions).
@@ 3 Results
3.1 Datasets Description and Data Pre-processing
We utilized three diverse datasets, comprising a total of 1,623 brain scans, to evaluate the effectiveness of our proposed methods. These datasets span various age groups, from 34 gestational weeks to elderly adults, and include both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD). Specifically, we used T1-weighted structural MRI scans from 1,064 young adults in the Human Connectome Project (HCP) S1200 release; 282 elderly adults, and 200 AD patients from the ADNI dataset; and 27/30/20 subjects in 2-year/6-month/31-week groups from the dHCP datasets. The pre-processing steps included brain skull removal, tissue segmentation, and cortical surface reconstruction conducted via the FreeSurfer package [FISCHL2012774].
@@ 3.2 Constructed GS-Net
In our experiments, we constructed GS-Net for each individual and presented the results of 25 randomly selected subjects from five different cohorts, including AD patients, elderly adults, young adults, individuals at 6 months, and individuals at 34 weeks gestation. These results are illustrated in Fig. 2. Each subject’s sphere surface is color-coded by gyri altitude, depicting the specific patterns of gyri and sulci of the individual. The generated GS-Net is overlaying on the sphere surface, comprising five components: GyralNet (blue lines), SulcalNet (red lines), edges connecting GyralNet and SulcalNet (yellow lines), GC (gyri conjunction) nodes (green bubbles), and SC (sulci conjunction) nodes (orange bubbles). Notably, the deepest red and blue regions on the sphere surface represent the locations with the highest positive and negative gyri altitude, corresponding to the primary gyri crests and sulci Valleys. Our generated GS-Net accurately covers these regions and effectively connects gyri regions and sulci regions along the folding pattern into a unified network. This highlights the precision and effectiveness of GS-Net.
@@ 3.3 Comparison with Baseline Methods
We compared our method with two baselines: a single-measure sulc-based approach and the method of [chen2017gyral]. Because cortical folding patterns are highly complex, relying on a single measure is more sensitive to noise and often leads to inaccurate 3HG localization. As shown in Fig. 3 , these baselines frequently produce redundant 3HG detections (yellow circles). Previous method also uses a single measure and depends on multiple hand-crafted stopping parameters during tree construction and pruning. Due to large inter-subject variability, one parameter set cannot generalize well, causing good performance in some subjects but redundant or missing 3HG branches in others. In contrast, our approach jointly uses sulc and curv, providing richer geometric cues, and employs an adaptive termination criterion based on complete erosion that naturally identifies the population-consistent gyral crest. This eliminates the need for fixed thresholds and yields more accurate and stable 3HG detection across individuals.
@@ 3.4 Runtime Comparison
We report the runtime for 30 randomly selected subjects, where the horizontal axis denotes subjects and the vertical axis indicates execution time (Fig. 4). The traditional method requires extremely long processing time, often taking several hours per subject because the large amount of remaining gyral regions leads to time-consuming tree construction and pruning. In contrast, our Multi-Measures approach achieves a dramatically shorter and far more stable runtime, typically completing the entire pipeline within only a few seconds. This efficiency highlights the scalability of our framework and its suitability for large-scale population studies.
@@ 4 Discussion and Conclusion
In this work, we introduced GS-Net, a framework that integrates the brain’s two primary folding patterns, gyri and sulci, into a unified anatomical network. Unlike traditional region-based approaches, GS-Net represents fine-scale anatomical landmarks as network nodes, enabling more detailed and biologically meaningful characterization of cortical morphology. We evaluated GS-Net on multiple datasets totaling over 1,600 brains across diverse cohorts and age groups, demonstrating that GS-Net accurately captures individual folding patterns. Compared with baseline methods, GS-Net not only achieves higher accuracy but also significantly reduces computation time, providing a fast and robust tool for large-scale brain network studies.

### v2

@@ Abstract
Our brain functions as a complex communication network, and studying it from a network perspective offers valuable insights into its organizational principles and links to cognitive functions and brain disorders. However, most current network studies typically use brain regions as nodes, often overlooking the intricate folding patterns of finer-scale anatomical landmarks within these regions. In this study, we introduce a novel approach to integrate the brain’s two primary folding patterns – gyri and sulci – into a unified network termed the Gyral-Sulcal-Net (GS-Net), in which three different types of finer-scale landmarks have been successfully identified. We evaluated the proposed GS-Net across multiple datasets, comprising over 1,600 brain scans, spanning different age groups (from 34 gestational weeks to elderly adults) and cohorts (healthy brains and those with pathological conditions). The experimental results demonstrate that the GS-Net can effectively integrate and represent diverse cortical folding patterns from a network perspective. More importantly, this approach offers a promising way for integrating different folding patterns into a unified anatomical brain network, alongside structural and functional networks, providing a comprehensive framework for studying brain networks.
@@ 1 Introduction
Brain network study starts with the identification of network nodes as interacting units and their interconnections as edges. In neuroimaging studies, various parcellation methods [ARSLAN20185], including those based on anatomical landmarks (e.g., AAL), cytoarchitectonic information (e.g., Brodmann areas), and connectivity patterns, are employed to subdivide the brain into a set of brain regions, which are utilized as network nodes. Advanced imaging techniques, such as Magnetic Resonance Imaging (MRI), are employed to construct macro-scale networks based on these ROIs, including the structural network inferred from diffusion MRI (dMRI) and the functional network estimated using resting-state functional MRI (rs-fMRI) [doi:10.1073/pnas.0811168106, https://doi.org/10.1002/hbm.22933]. Structural network allows estimation of the physical connections, while the functional network elucidates putative functional connections. Brain network studies have made significant breakthroughs in understanding the fundamental organizational principles of the normal brain [ZHANG2022102463] and brain diseases [Bassett9239]. However, current research still lacks sufficient investigation into two critical aspects. Firstly, most existing brain networks utilize brain regions as nodes, which often cover large areas at a relatively coarse scale. The folding pattern of finer-grained anatomical landmarks within these brain regions tends to be overlooked, such as sulci fundi (deepest loci of sulci) and gyri peaks (highest loci in gyri) [Thecerebralsulciandgyri]. Accumulating evidences suggest that these landmarks confer specific structural, functional, and cognitive patterns [10.1093/cercor/bhac537] and are under a stronger genetic or heritability influence than other cortical regions [10.1093/cercor/bhac537, IM2019881]. Therefore, establishing a finer-scale brain network based on these landmarks is significant and would serve as a valuable complement to the region-based brain network. Secondly, existing studies primarily employ structural and functional connectivity to construct brain networks, and anatomical folding patterns cannot be effectively integrated into such frameworks. Although certain studies have investigated the unique structural-functional patterns of specific folding patterns and their changes in disease development [KIKINIS19947, 10.3389/fpsyt.2022.1033918], these methods often independently analyze particular folding patterns. They fail to effectively integrate different folding patterns within a unified brain network, alongside the brain structural and functional networks, to comprehensively study the brain from a network perspective. To address these problems, in this study, we introduced a novel framework to integrate the brain’s two primary folding patterns, gyri and sulci, into a unified network called Gyral-Sulcal-Net (GS-Net), while also automatically identifying three important landmarks – gyri conjunctions, sulci conjunctions, and sulci-gyri conjunctions. We evaluated the proposed GS-Net using multiple datasets comprising 1,600+ brain scans. These datasets cover various age groups ranging from 34 gestational weeks to elderly adults, and different cohorts including healthy brains and brains with pathological conditions. The generated GS-Net can accurately describe the folding pattern of the brain from a network perspective.
@@ 2.1 Method Overview
Fig. 1: The pipeline of the proposed GS-Net framework.
@@ 2.4 Tree Marching and Trimming
After segmenting the gyri crests and sulci valleys, we employed tree marching and trimming methods (Fig. 1 (c)) to construct networks upon gyri crests and sulci valleys. Taking the gyri crest regions as an example, the tree marching process begins with the vertex at the highest gyri altitude and progressively connects other vertices in descending order until all vertices within the gyri crests are linked. During this process, all vertices in the crest area are connected, resulting in some redundant branches extending from the skeleton to the boundary. We then perform trimming to eliminate these extraneous branches. Since the remaining gyri crests form narrow regions, the redundant branches that extend beyond the main structure are relatively short and can be easily identified. Following the trimming, the main skeleton is integrated into a network that delineates the primary morphology of the gyri regions, which we refer to as GyralNet. Similarly, SulcalNet is constructed for the sulci valley regions.
@@ 2.5 Integration and Landmark Identification
GS-Net contains three distinct types of nodes, each playing a key role in brain networks. The first type includes conjunctions of multiple edges in GyralNet, representing intersections of various gyri on the cerebral cortex. These nodes have been highlighted in the literature as regions with thicker cortices, higher fiber density, and significant variations in structural and functional connectivity [LI20101202, 10.1093/cercor/bhx227]. Based on their connections to free sulci, we further classify these nodes as GC (Gyri Conjunction) and SGC (Sulci-Gyri Conjunction) in this study. Additionally, the conjunctions of multiple edges in SulcalNet show unique structural, functional, and genetic characteristics. Existing study suggests these nodes exhibit spatial consistency across individuals during development [Thecerebralsulciandgyri] and are subject to stronger genetic or heritability influences compared to other cortical regions [10.1093/cercor/bhac537]. These nodes are referred to as SC (Sulci Conjunction).
@@ 3 Results
3.1 Dataset Description and Data Pre-processing
We utilized three diverse datasets, comprising a total of 1,623 brain scans, to evaluate the effectiveness of our proposed methods. These datasets span various age groups, from 34 gestational weeks to elderly adults, and include both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD). Specifically, we used T1-weighted structural MRI scans from 1,064 young adults in the Human Connectome Project (HCP) S1200 release; 282 elderly adults, and 200 AD patients from the ADNI dataset; and 27/30/20 subjects in 2-year/6-month/34-week groups from the dHCP datasets. The pre-processing steps included brain skull removal, tissue segmentation, and cortical surface reconstruction conducted via the FreeSurfer package [FISCHL2012774].
@@ 3.2 Constructed GS-Net
In our experiments, we constructed GS-Net for each individual and presented the results of 25 randomly selected subjects from five different cohorts, including AD patients, elderly adults, young adults, individuals at 6 months, and individuals at 34 weeks of gestation. These results are illustrated in Fig. 2. Each subject’s sphere surface is color-coded by gyri altitude, depicting the specific patterns of gyri and sulci of the individual. The generated GS-Net is overlaying on the sphere surface, comprising five components: GyralNet (blue lines), SulcalNet (red lines), edges connecting GyralNet and SulcalNet (yellow lines), GC (gyri conjunction) nodes (green bubbles), and SC (sulci conjunction) nodes (orange bubbles). Notably, the deepest red and blue regions on the sphere surface represent the locations with the highest positive and negative gyri altitude, corresponding to the primary gyri crests and sulci valleys. Our generated GS-Net accurately covers these regions and effectively connects gyri regions and sulci regions along the folding pattern into a unified network. This highlights the precision and effectiveness of GS-Net.
@@ 3.3 Comparison with Baseline Methods
We compared our method with two baselines: a single-measure sulc-based approach and the method of [chen2017gyral]. Because cortical folding patterns are highly complex, relying on a single measure is more sensitive to noise and often leads to inaccurate 3-hinge gyrus (3HG) localization. As shown in Fig. 3, these baselines frequently produce redundant 3HG detections (yellow circles). The previous method [chen2017gyral] also uses a single measure and depends on multiple hand-crafted stopping parameters during tree construction and pruning. Due to large inter-subject variability, one parameter set cannot generalize well, causing good performance in some subjects but redundant or missing 3HG branches in others. In contrast, our approach jointly uses sulc and curv, providing richer geometric cues, and employs an adaptive termination criterion based on complete erosion that naturally identifies the population-consistent gyral crest. This eliminates the need for fixed thresholds and yields more accurate and stable 3HG detection across individuals.
@@ 3.4 Runtime Comparison
We report the runtime for 30 randomly selected subjects, where the horizontal axis denotes subjects and the vertical axis indicates execution time (Fig. 4). The traditional method requires the extremely long processing time, often taking several hours per subject, because the large amount of remaining gyral regions leads to time-consuming tree construction and pruning. In contrast, our Multi-Measures approach achieves a dramatically shorter and far more stable runtime, typically completing the entire pipeline within only a few seconds. This efficiency highlights the scalability of our framework and its suitability for large-scale population studies.
@@ 4 Discussion and Conclusion
In this work, we introduced GS-Net, a framework that integrates the brain’s two primary folding patterns, gyri and sulci, into a unified anatomical network. Unlike traditional region-based approaches, GS-Net represents fine-scale anatomical landmarks as network nodes, enabling more detailed and biologically meaningful characterization of cortical morphology. We evaluated GS-Net on multiple datasets totaling over 1,600 brain scans across diverse cohorts and age groups, demonstrating that GS-Net accurately captures individual folding patterns. Compared with baseline methods, GS-Net not only achieves higher accuracy but also significantly reduces computation time, providing a fast and robust tool for large-scale brain network studies.